the measure of angle X is 12 less than twice the measure of angle Y. If angle X and angle Y are supplementary angles, find the measure of angle X
Answer by checkley77(12844)

X=2Y+12
X+Y=180
X=180-Y
2Y+12=180-Y
2Y+Y=180-12
3Y=168
Y=168/3
Y=56 ANS.
X+56=180
X=180-56
X=124 ANS.
PROOF:
124=2*56+12
124=112+12
124=124
56+124=180
180=180